This year's soybean crop of 224,000 tons represents an increase of 40% over last year's crop. How many tons of soybean were produced last year?
Answer by nyc_function(2741)

224,000 * 0.40 = 89,600
soybeans this year = 224,000
last year = 224,000 - 89,600
224,000 - 89,600 = 134,400
There were 134,400 soybeans last year.